Resource Consider the information about the economy of Pakistan. Note that the currency of Pakistan is the rupee. The government purchases: 2.50 trillions of rupees. Individuals consume: 8.80 trillions of rupees. Individuals save:4.97 trillions of rupees. Businesses invest: 1.20 trillions of rupees. Foreigners spend: 0.56 trillions of Rupees to purchase Pakistani firms. Pakistan imports: 2.17 trillions of rupees. Pakistan exports: 1.27 trillions of rupees. Calculate Pakistans GDP. Assume that the values are all current and no conversions need to be made. Pakistans GDP: 11.99 trillions of rupees
Answer #1

To calculate: Pakistan's GDP

Formula: GDP = C + I + G + X - M

Given:

Government purchases (G) = 2.50 trillions

Individual consumption (C) = 8.80 trillions

Businesses invest (I) = 1.20 trillions

Pakistan imports (M) = 2.17 trillions

Pakistan exports (X) = 1.27 trillions

(Individual savings and foreigners spending are distractors - Not required to calculate GDP)

Calculation:

GDP = C + I + G + X - M

(Substitute the given values in the formula)

GDP = 8.80+1.20+2.50+1.27-2.17 = 11.6

Pakistan's GDP = 11.6 trillions of rupees